OpenStudy (anonymous):

A gym offers regular memberships for $80 per month and off-peak memberships for $60 per month. Last month, the gym sold a total of 420 memberships for a total of $31,100. How many memberships sold were regular memberships? A.)125 B.) 140 C.) 235 D.) 295

OpenStudy (whpalmer4):

let \(r\) be the number of regular memberships. If they sold 420 memberships, then \(420-r\) is the number of off-peak memberships. Revenue from regular memberships at $80 each is \(80r\) Revenue from off-peak memberships at $60 each is \(60(420-r)\) Total revenue is regular + off-peak = $31,100 \[31,100 = 80r + 60(420-r)\] Solve for \(r\) and you've got your regular membership count.

OpenStudy (anonymous):

So is it A

OpenStudy (anonymous):

That is what i got

OpenStudy (whpalmer4):

If 125 regular memberships were sold, how many off-peak memberships were sold?

OpenStudy (anonymous):

295

OpenStudy (anonymous):

So is the answer D

OpenStudy (whpalmer4):

Stop guessing. You can work the problem out from what I gave you, and KNOW that you have the right answer. Here's how to check your answer. Suppose you think the answer is A, 125 regular memberships (which implies 295 off-peak memberships). 125 * 80 + 295*60 = 10000 + 17700 = 27700 27700 does not equal 31100, so A is not the correct answer.

OpenStudy (whpalmer4):

Can you show me your work that led you to think A was the correct answer? If you honestly did the problem and concluded that A was the correct answer, you must have made a mistake somewhere, and we should find the mistake and correct it so that you don't make similar mistakes in the future.